OpenStudy (anonymous):

Given the equation 5x + y = 7, which equation below would cause a consistent-independent system? 10x + 2y = 14 -15x - 3y = -6 5x + y = -7 6x + 2y = 7

OpenStudy (danjs):

independent, they will both have different slopes, meaning they will intersect once on the graph, and have one solution

OpenStudy (anonymous):

Danj, i'm so sorry to ask you this, but I have no clue what i'm doing when I see this question...

OpenStudy (danjs):

10x + 2y = 14 is the same line as 5x + y =7 Multiply the equation by 2, it is the same line, not independent

OpenStudy (anonymous):

Ahh

OpenStudy (danjs):

-15x - 3y = -6 the x and y coefficients -15 and -3, are -3 times the x and y coeffiecents in the original line. the two will be PARALLEL lines. No Solution, and Inconsistent

OpenStudy (danjs):

5x + y = -7 same as the last one, the coefficients on the X and Y are a multiple of the original line, this time just 1 times the original. Parallel Lines, No Solutions, Inconsistent

OpenStudy (anonymous):

So the last one is the answer?

OpenStudy (danjs):

Right, there is not number you can multiply 6x+2y=7 by to get the first equations coefficients on X and Y, If you solved these two equations, you will get one unique answer, the intersection point of the two lines (x,y) This is a consistent, independent, one solution, system.

OpenStudy (anonymous):

Thank you!

OpenStudy (danjs):

review that linked page in the first response, it will give you all that info again.